OpenStudy (anonymous):

what are the equations for the conservation of momentum for Inelastic and elastic collisions?

OpenStudy (jamesj):

A collision is elastic or not depending on whether KE is conserved or not, as we've seen in a few of your questions now. That being the case, Conservation of Momentum is unchanged whether KE is conserved or not.

OpenStudy (anonymous):

How do I find the change in Kinetic energy

OpenStudy (jamesj):

If a body of mass m has velocity v before an event and velocity v' after the event, then the change in KE is \[ KE_{after} - KE_{before} \] \[ = \frac{1}{2}mv'^2 - \frac{1}{2}mv^2 \] \[ = \frac{1}{2}m (v'^2 - v^2) \]